We have a rectangle where 2x+2y=45, and x=2/3y, solve for x (or Y).
How do you solve this?

Found 2 solutions by ichudov, dreamgirl9999:
Answer by ichudov(507)

it is a regular linear system, I have a solver for it.

Answer by dreamgirl9999(1)

First you replace the x in 2x+2y=45 with 2/3y since x=2/3y.
After doing this you should get 2(2/3y)+2y=45.
Now distribute and you should get 1_1/3y (or 4/3) +2y=45.
After combining the like terms: 3_1/3y=45
Now do the inverse operations by dividing by 3_1/3 on both sides to get y by itself:: y=13_1/2
Answers are y=13_1/2 (or 13.5) and x=9 (you get 9 by multiplying 2/3 by 13.5).